OpenStudy (anonymous):

What is the equation of the line that passes through the point (4, 1) and is perpendicular to the line y = –2x + 4? What is the equation of the line that passes through the point (4, 1) and is perpendicular to the line y = –2x + 4? @Mathematics

OpenStudy (anonymous):

\[(y - 1) = 2(x-4)\]\[y = 2x-7\]

OpenStudy (anonymous):

(y - 1) = 1/2(x - 4) y - 1 = 1/2x - 2 y = 1/2x - 1 Perpendicular means negative reciprocal slope

OpenStudy (anonymous):

ah yea! ur right i forgot the reciprocal